Calculate the freezing point depression of the given electrolytic solution: 5.7 g CaCl2 (calcium chloride) in 18.6 kg H2O (water)
1
Expert's answer
2020-10-28T08:40:19-0400
"\Delta T = iKC_m"

"C_m = {n(CaCl_2) \over m(H_2O)} = {m(CaCl_2) \over M(CaCl_2)*m(H_2O)}"

"\Delta T= 3*1.86*{5.7 \over 111*18.6} = 0.015"

"T_{freezing} = 0^oC-0.015^oC = -0.015^oC"